find three consecutive positive integers such that the product of the first and third, minus the second, is 1 more than 6 times the third.
What is the smallest integer?
What is the second smallest integer?
What is the largest integer?

Answer by ewatrrr(24785)

 
Hi,
find three consecutive positive integers
What is the smallest integer? n
What is the second smallest integer? (n+1)
What is the largest integer? (n+2)
product of the first and third, minus the second, is 1 more than 6 times the third.
n(n+2) - (n+1) = 6(n+2)+1
n^2 -5n -14 = 0
(n+2)(n-7) = 0, +n+=+7 (tossing out the negative solution for n)
the three consecutive positive integers are: 7,8,9
CHECKING our Answer***
7*9 - 8 = 55 = 6*9 + 1